• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name M Friend

    Log entry time 02:56:37 on October 5,2009

    Entry number 293194

    keyword=Compton Run Summary

    We attempted to run in electron singles mode at 0.0mm, but the beam halo is clearly seen for 5 strips, so we cannot run like this right now.

    Electron detector moved back to 1.0mm, still in electron singles mode, still triggering FADC on electron detector, 3 plane coincidence.

    00:11 Run 20088 started, 3 plane electron only coincidences. There was some fast feedback problems at the begining of this run, and the beam position was changing a lot, but rates looked okay on the strip chart the whole time. Run ended at 00:28.

    Switch to 1 plane coincidence.

    00:29 Start run 20089, 1 plane electron only run, FADC trig on edet. Ended at 00:44.

    FADC trigger set back to PMT trigger, standard DAQ set to photons + coincidences e-g, 3 plane electron coincidences, prescales set to 4% for photon, 0% for electron, 100% for coincidences.

    Electron detector moved to 0.0mm for coincidence run, per Alex's request.

    00:54 Start run 20091, e-g coincidences at 100uA. Run ended at 1:05 so that it can be analyzed and we can make sure there's data before we continue. Data is there, but halo means that we shouldn't run like this.

    Electron detector moved back to 1.0mm, had to reboot motion ioc (again).

    1:35 Start run 20092, still e-g coincidences at 100uA, this time at edet position 1.0mm, still with 3 plane electron coincidences. Ended at 2:14.

    Edet moved to 0.0mm, FADC trigger changed to edet, spyconf changed to electron only, still with 3 plane coincidences.

    Requested 5uA beam by closing the slits. There was some problem with the first run, so we requested 100uA again to check the ETROC. However the ETROC looks okay. The new plan:

    -Take a production run for two hours
    -Check an electron only run at 100uA at 0.0mm
    -Go to 5uA
    -Take a short electron only run at 0.0mm at 5uA to check beam halo
    -Take a 40 min coincidence run at 5uA, at 0.0mm if possible, at 1.0mm otherwise. Then back up to 100uA.
    -Production for the rest of the night.

    Edet motion ioc needed to be rebooted, edet moved to garage.

    Spyconf run mode changed to photon only, FADC config chaged to standard run.

    2:42 Start run 20095, standard production run.